1 inch corresponds to 5 feet find the length of a wall represented by a wall 6and 2/3
Answer by checkley79(3341)

1/5=6.67/x
x=5*6.66
x=33.3 ans.
OR
1/5=x/6.67
5x=6.67
x=6.67/5
x=1.33 ans.